# When do Neural Networks Learn World Models?

### v2

@@ Abstract
Humans develop world models that capture the underlying generation process of data. Whether neural networks can learn similar world models remains an open problem. In this work, we provide the first theoretical results for this problem, showing that in a multi-task setting, models with a low-degree bias provably recover latent data-generating variables under mild assumptions—even if proxy tasks involve complex, non-linear functions of the latents. However, such recovery is also sensitive to model architecture. Our analysis leverages Boolean models of task solutions via the Fourier-Walsh transform and introduces new techniques for analyzing invertible Boolean transforms, which may be of independent interest. We illustrate the algorithmic implications of our results and connect them to related research areas, including self-supervised learning, out-of-distribution generalization, and the linear representation hypothesis in large language models.
@@ 1 Introduction
    In Section 5, we illustrate the algorithmic implications of our results on two representative tasks: polynomial extrapolation (Xu et al., 2021) and learning physical laws (Kang et al., 2024). We show that architectures inspired by our analysis outperform conventional architectures such as ReLU MLPs and transformers (Vaswani et al., 2017) in these tasks.
@@ 2 Formulation of Learning World Models
This definition resembles the data generation process used in many latent variable models such as non-linear ICA (Hyvärinen & Pajunen, 1999), invariant feature learning (Arjovsky et al., 2019), and causal representation learning (Schölkopf et al., 2021). The main difference is that unlike these models, we do not assume $p({\mathbf{z}})$ to be any structured distribution. Intuitively, a natural way to formalize “understanding” this generation process is to approximately invert it, i.e., recover the latent variables ${\mathbf{z}}$ from the observed data ${\mathbf{x}}$. This gives our basic formulation of learning world models.
@@
[…]
$\qquad\forall{\bm{z}}\in\mathrm{supp}(p),{\bm{x}}=\psi({\bm{z}})\}.$
If a function composition $h_{(1)}\circ\cdots\circ h_{(q)}$ is in ${\mathcal{H}}(h)$, then we say it is a realization of $h$.
@@
To address this puzzle, in this work we propose to incorporate the implicit bias of neural networks instead of using task performance as the only identifiability criterion. Specifically, we explore if the bias towards low-complexity realizations, a trait of both human reasoning and deep learning, could be used to steer the realization towards non-trivial recovery of the true latents. Yet, the main challenge is that for continuous functions with inputs in ${\mathbb{R}}^{d}$, we lack a well-established complexity measure that is amenable to analysis. For example, Kolmogorov complexity (Li et al., 2008) in algorithmic learning theory offers a unified framework for defining the complexity of any object, yet it is uncomputable. Fortunately, we will next show that this problem could be circumvented by using Boolean models.
@@ 3 Complexity Measures
A direct consequence of the Boolean modeling of variables is that all functions involved are Boolean functions (see Section B for a basic introduction). Notably, this itself does not resolve the non-identifiability issue present in the continuous case, as functions with Boolean inputs and outputs can still exhibit arbitrary nonlinearity. Instead, the primary advantage of this approach lies in the useful machinery it offers for defining functional complexity through the Fourier-Walsh transform:
@@
The Fourier-Walsh transform shows that every Boolean function can be represented as a linear combination of parity functions $\chi_{S}$ that capture all non-linear relationship between inputs and outputs. In fact, it can be shown that parity functions are a basis of the vector space ${\mathcal{F}}^{n}\vcentcolon=\{f:\{\pm 1\}^{n}\to{\mathbb{R}}\}$ of $n$-dimensional Boolean functions (see Section B for more details). By contrast, the space of arbitrary continuous functions does not have a similar, theory-friendly basis.
@@
Note that the degree of $f$ equals to the maximum degree of the basis functions it uses: $\mathrm{deg}(f)=\max\{\mathrm{deg}(\chi_{S}):\hat{f}(S)\neq 0\}$. This interpretation will be useful in Section 4.4.
@@
For a realization $h_{(1)}\circ\cdots\circ h_{(q)}$, its realization degree is
$\widehat{\mathrm{deg}}(h_{(1)}\circ\cdots\circ h_{(q)})=\sum_{i\in[q]}\mathrm{deg}(h_{(i)}).$ (4)
For example, the realization degree of a flat realization $h^{*}\in{\mathcal{H}}(h)$ coincides with its degree; the realization degree of a hierarchical realization $h=g\circ\Phi$ is $\widehat{\mathrm{deg}}(g\circ\Phi)=\mathrm{deg}(g)+\mathrm{deg}(\Phi)$. Compared to degree, realization degree better reflects the cost of implementing each function in hierarchical realizations as in practice. Throughout this work, we say a model exhibits a low-degree bias if it minimizes the realization degree. We conclude this section by two definitions that will be useful in characterizing the impact of the low-degree bias on flat and hierarchical realizations.
@@
For a task $h$ and representation $\Phi:{\mathcal{X}}\to{\mathcal{Z}}$, let $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$, and the conditional degree of $h$ on the representation $\Phi$ is defined as
@@
By Definition 3.5, flat realizations with the low-degree bias satisfy $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$. Analyzing the impact of representations is more involved; Definition 3.6 suggests that a representation $\Phi$ only makes a task $h$ “simpler” if $\mathrm{deg}(h\mid\Phi)>0$, i.e., solving the task on top of $\Phi$ has a smaller degree compared to low-degree flat realizations. We will explore this notion further in the next section.
@@ 4 Theoretical Analysis
This section presents our main theoretical results. We first study how the low-degree bias introduced in Section 3 drives representation learning in a multi-task setting (Section 4.1). We then present our main results on learning world models, showing sufficient conditions for the identification of latent data-generating variables (Section 4.2). Next, we show provable benefits of learning world models (Section 4.3). We conclude this section by a study on how the model architecture impacts world model learning (Section 4.4). All proofs are deferred to Section C.
@@ 4.2 Conditions for Learning World Models
We begin by characterizing the space of all possible tasks. Given that the observed data is generated by ${\mathbf{x}}=f({\mathbf{z}})$ with an invertible $f$, every task $h:{\mathcal{X}}\to{\mathbb{R}}$ can be equivalently defined via a function $h^{\prime}:{\mathcal{Z}}\to{\mathbb{R}}$ as $h=h^{\prime}\circ\psi^{-1}$. Thus, the space of all possible tasks on ${\mathcal{X}}$ is represented as
@@
Theorem 4.4 can be viewed as a “no free lunch”-like theorem for representation learning. The original no free lunch theorem (Wolpert, 1996) states that every learner’s performance is equally good when averaged over a uniform distribution on learning problems; here we show that every viable representation is equally complex when averaged over a uniform distribution on tasks. As we will show in Section C.4, the technical intuition of this result is that every representation renders some tasks in the task space “simple” and others “complex”, with the overall task-averaged complexity independent of the particular choice of the representation.
@@
  • A limitation of Theorem 4.8 is that we requires a non-zero probability of explicitly sampling degree-$1$ tasks, in which latent variables are also task outputs. However, we emphasize that this probability is exponentially small as $d$ becomes large, and we conjecture that it can be completely removed in many settings. See Section C.6 for more discussion.
  • Technically, in our proof we introduce a degree analysis of $k$-degree Boolean functions composed with invertible transforms (Lemma C.10), which may be of independent interest in analyzing similar problems.
Connection to the linear representation hypothesis. A number of recent mechanistic interpretation studies show that LLMs often represent abstract, interpretable features as directions in their intermediate representation space (Nanda et al., 2023; Marks & Tegmark, 2023; Gurnee & Tegmark, 2024). Theorem 4.8 can be viewed as a provable, Boolean version of the emergence of such linear representations: permutations and negations are precisely all degree-$1$ Boolean functions, a natural counterpart of degree-$1$ real polynomials (i.e., linear functions) in the real domain. The main significance of this result is that even if proxy tasks can be complex, non-linear functions over true latent variables ${\bm{z}}$, we can still recover ${\bm{z}}$ up to very simple transforms.
Connection to the linear representation hypothesis. A number of recent mechanistic interpretation studies show that LLMs often represent abstract, interpretable features as directions in their intermediate representation space (Nanda et al., 2023; Marks & Tegmark, 2023; Gurnee & Tegmark, 2024). Theorem 4.8 can be viewed as a provable, Boolean version of the emergence of such linear representations: permutations and negations are precisely all degree-$1$ Boolean functions, a natural counterpart of degree-$1$ real polynomials (i.e., linear functions) in the real domain. The main significance of this result is that even when proxy tasks involve complex, non-linear functions over true latent variables ${\bm{z}}$, we can still recover ${\bm{z}}$ up to very simple transforms despite the presence of such nonlinearity.
@@ 4.3 Benefits of Learning World Models
Up to now, we have presented sufficient conditions for learning world models. As a complement of these results, this section demonstrates provable benefits of learning world models in the context of an out-of-distribution generalization setting introduced by Abbe et al. (2023).
@@
Let the latent variables during training be uniformly sampled from the Hamming ball $B_{r}\vcentcolon=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$ with $r<d$, and let those during testing be uniformly sampled from ${\mathcal{Z}}$. Let $h:\{\pm 1\}^{m}\to\{\pm 1\}$ be a downstream task such that $h\circ\psi$ is a parity function with degree $q\geq k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose i}\rceil$. Then, if $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$, the following hold: (i) the test mean square error (MSE) of any $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$ is larger than $1$; (ii) let $\Phi^{*}$ be a representation that learns the world model up to negations and permutations as in Theorem 4.8 and let $g^{*}$ be a function such that $g^{*}\circ\Phi^{*}\in{\mathcal{H}}(h)$, then the test MSE of $g^{*}\circ\Phi^{*}$ is 00.
@@
As also noted by Abbe et al. (2023), a practical scenario reflected by sampling from $B_{r}$ is length generalization of transformers (Anil et al., 2022; Press et al., 2022). Here we show that in this setting, a hierarchical realization with the world model is provably more generalizable than any flat realization, despite that both of them achieve zero i.i.d. test error.
@@ 4.4 Impact of Model Architecture
In the above analysis, we study the role of the low-degree bias with the assumption that the task solutions perfectly adhere to it in the function space. Yet, practical training of neural networks is often more involved than this abstraction. Although neural networks are known as universal function approximators (Hornik et al., 1989; Funahashi, 1989), prior work has shown that models with different architectures may represent the same function differently (Raghu et al., 2021). In particular, embedded nonlinearities such as activation functions can steer how functions are represented by neural networks (Xu et al., 2020; Ziyin et al., 2020; Teney et al., 2024).
@@
Notably, for a given input dimension $n$, the Fourier-Walsh transform uses parity functions $\chi_{S}$ as a basis of the $2^{n}$-dimensional vector space ${\mathcal{F}}^{n}=\{f:\{\pm 1\}\to{\mathbb{R}}\}$ (see Definition 3.1). To obtain a different basis, we define a basis transform $U$, i.e., an invertible linear transform on ${\mathcal{F}}^{n}$ such that $\{U(\chi_{S})\mid S\subseteq[n]\}$ is a basis of ${\mathcal{F}}^{n}$. The degree of a function $f:\{\pm 1\}^{n}\to{\mathbb{R}}$ under the new basis is then given by
@@ 5 Algorithmic Implications
In this section, we illustrate the algorithmic implications of basis compatibility through two representative tasks: polynomial extrapolation (Xu et al., 2021) and learning physical laws (Kang et al., 2024; Motamed et al., 2025). The first task provides an illustrative example of our results in Section 4.4, while the second is of greater practical interest for learning real-world models and is related to recent efforts in developing video prediction models as world simulators (Brooks et al., 2024). While our experiments are currently limited in scale and mainly serve as proof-of-concept demonstrations, we view extending the ideas discussed in this section to broader contexts as a promising future direction.
@@ 5.1 Polynomial Extrapolation
We begin by a synthetic task in which we train multilayer perceptrons (MLPs) to fit real polynomials $P_{n}(x)=\sum_{i=0}^{n}a_{i}x^{n}$ (see Section D.1 for details). The space of polynomials has a standard basis $\{1,x,x^{2},\ldots\}$, making it a natural counterpart of the Boolean function space with the parity basis. Recovering all basis functions used by $P_{n}$ could generate any polynomial by linearly combining these basis functions, which enables extrapolation.
@@
Motivated by this explanation, here we provide a simple fix: replacing a portion of the ReLU activation functions with functions that are more compatible with the polynomial basis. In practice, we replace half of ReLU functions in each MLP layer by one of the two functions including the identity function $\sigma(x)=x$ and the quadratic function $\sigma(x)=x^{2}$. Representing $f(x)=x^{k}$ would be much easiler using these functions, and we thus expect that a neural network with the low-complexity bias would then use them instead of the remaining ReLU despite the same expressive ability they have, resulting in $\hat{f}\approx f$. As shown in Figure 2(b), this simple method indeed leads to significant improvement in extrapolation. Please see Section E.1 for more results and discussion.
@@
We now show that the approach in Section 5.1 also benefits a sequence prediction task aiming at learning physical laws. Correctly abstracting fundamental physical laws is essential for any model to be a real “world model” for the physical world (Motamed et al., 2025). Notably, a model that “understands” the physical laws is expected to generalize these laws to unseen distributions rather than fit them only in the training domain. Following Kang et al. (2024), we generate two types of object motion sequences reflecting basic physical laws: (i) single-object parabolic motion, and (ii) two-object elastic collision motion (see Section D.2 for more details). We train a transformer (Vaswani et al., 2017) to predict the motion of objects conditional on the first few frames in the motion sequence. The model is then evaluated in an out-of-distribution generalization setup with objects having different initial velocities and sizes. To apply our method, we simply replace every MLP in the transformer with our modified MLP in Section 5.1 (we also replace the remaining ReLUs with GELUs).
As shown in Figure 2(c), our model achieves lower prediction error than transformer in both settings. Note that both models achieve near zero training error. Thus, the fact that our model generalizes better out-of-distribution is not because it fits the data better, but due to it better capturing the underlying laws in object movements. See Section E.2 for the predicted motions from both models.

### v3

@@ Abstract
Humans develop world models that capture the underlying generation process of data. Whether neural networks can learn similar world models remains an open problem. In this work, we present the first theoretical results for this problem, showing that in a multi-task setting, models with a low-degree bias provably recover latent data-generating variables under mild assumptions—even if proxy tasks involve complex, non-linear functions of the latents. However, such recovery is sensitive to model architecture. Our analysis leverages Boolean models of task solutions via the Fourier-Walsh transform and introduces new techniques for analyzing invertible Boolean transforms, which may be of independent interest. We illustrate the algorithmic implications of our results and connect them to related research areas, including self-supervised learning, out-of-distribution generalization, and the linear representation hypothesis in large language models.
@@ 1 Introduction
    In Section 5, we illustrate the algorithmic implications of our results on two representative tasks: polynomial extrapolation (Xu et al., 2021) and learning physical laws (Kang et al., 2025). We show that architectures inspired by our analysis outperform conventional architectures such as ReLU MLPs and transformers (Vaswani et al., 2017) in these tasks.
@@ 2 Formulation of Learning World Models
This definition resembles the data generation process used in many latent variable models such as non-linear ICA (Hyvärinen & Pajunen, 1999), invariant feature learning (Arjovsky et al., 2019), and causal representation learning (Schölkopf et al., 2021). The main difference is that unlike these models, here we do not assume $p({\mathbf{z}})$ to be any structured distribution.
Given Definition 2.1, a natural way to formalize “understanding” the generation process of data is to invert it, i.e., recover the latent variables ${\mathbf{z}}$ from the observed data ${\mathbf{x}}$. This leads to our basic formulation of learning world models.
@@
${\mathcal{H}}(h)\vcentcolon=\{f:{\mathcal{X}}\to{\mathbb{R}}\mid$ $f({\bm{x}})=h({\bm{x}}),$ (1)
$\forall{\bm{z}}\in\mathrm{supp}(p),{\bm{x}}=\psi({\bm{z}})\}.$
If a function composition $f_{(1)}\circ\cdots\circ f_{(q)}$ is in ${\mathcal{H}}(h)$, then we say it is a realization of $h$.
@@
To address this puzzle, in this work we propose to incorporate the implicit bias of neural networks into analysis instead of using task performance as the only identifiability criterion. Specifically, we explore if the bias towards low-complexity realizations, a trait of both human reasoning and deep learning models, could be used to steer the learned realization towards non-trivial recovery of true latents. Yet, the main challenge is that for continuous functions with inputs in ${\mathbb{R}}^{d}$, we lack a well-established complexity measure that is amenable to analysis. For example, Kolmogorov complexity (Li et al., 2008) in algorithmic learning theory offers a unified framework for defining the complexity of any object, yet being uncomputable. Fortunately, in the following sections we will show that this problem could be circumvented by using Boolean functions.
@@ 3 Complexity Measures
A direct consequence of the Boolean modeling of variables is that all functions involved are Boolean functions. Notably, this itself does not resolve the non-identifiability issue present in the continuous case, as functions with Boolean inputs and outputs can still exhibit arbitrary nonlinearity. Instead, the primary advantage of this approach lies in the useful machinery it offers for defining functional complexity through the Fourier-Walsh transform:
@@
The Fourier-Walsh transform shows that every Boolean function can be represented as a linear combination of parity functions $\chi_{S}$ that capture all non-linear relationship between inputs and outputs. In fact, it can be shown that parity functions are a basis of the vector space ${\mathcal{F}}^{n}\vcentcolon=\{f:\{\pm 1\}^{n}\to{\mathbb{R}}\}$ of $n$-dimensional Boolean functions (see Section D for more details). By contrast, the space of arbitrary continuous functions does not have a similar, theory-friendly basis. See Section B for more discussion on the use of Boolean functions and complexity measures.
@@
Note that the degree of $f$ equals to the maximum degree of the basis functions it uses: $\mathrm{deg}(f)=\max\{\mathrm{deg}(\chi_{S}):\hat{f}(S)\neq 0\}$. This interpretation will be useful when we consider different basis functions as in Section 4.4.
@@
Let $h$ be a task. For a realization $f_{(1)}\circ\cdots\circ f_{(q)}$ of $h$, its realization degree is defined as
$\widehat{\mathrm{deg}}(f_{(1)}\circ\cdots\circ f_{(q)})=\sum_{i\in[q]}\mathrm{deg}(f_{(i)}).$ (4)
For example, the realization degree of a flat realization $h^{*}\in{\mathcal{H}}(h)$ coincides with its degree; the realization degree of a hierarchical realization $h=g\circ\Phi$ is $\widehat{\mathrm{deg}}(g\circ\Phi)=\mathrm{deg}(g)+\mathrm{deg}(\Phi)$. Compared to degree, realization degree takes into account the cost of implementing each function in hierarchical realizations, which better captures representation learning in practice. Throughout this work, we say a model exhibits a low-degree bias if it minimizes the realization degree.
Finally, we introduce two definitions that are useful in characterizing the impact of the low-degree bias on flat and hierarchical realizations.
@@
For a task $h$ and a representation $\Phi:{\mathcal{X}}\to{\mathcal{Z}}$, let $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$ and assume that the set $\{g:{\mathcal{Z}}\to{\mathbb{R}}\mid g\circ\Phi\in{\mathcal{H}}(h)\}$ is not empty. The conditional degree of the task $h$ on the representation $\Phi$ is then defined as
@@
By Definition 3.5, flat realizations with the low-degree bias satisfy $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$. Analyzing the impact of representations is more involved; the main intuition behind Definition 3.6 is that a representation $\Phi$ only makes a task $h$ “simpler” if the conditional degree $\mathrm{deg}(h\mid\Phi)>0$, i.e., solving the task on top of $\Phi$ yields a smaller degree compared to low-degree flat realizations. This notion will be explored further in the next section.
@@ 4 Theoretical Analysis
This section presents our main theoretical results. We first study how the low-degree bias introduced in Section 3 drives representation learning (Section 4.1). We then present our main results on learning world models, showing sufficient conditions for identifying latent data-generating variables (Section 4.2). Next, we show provable benefits of learning world models in an out-of-distribution generalization setting (Section 4.3). We conclude this section by a preliminary analysis on how the model architecture impacts world model learning (Section 4.4). All proofs are deferred to Section E. Finally, Section C presents numerical experiments that substantiate our theoretical results.
@@ 4.2 Conditions for Learning World Models
We begin by characterizing the space of all possible tasks. Given that the observed data is generated by ${\mathbf{x}}=\psi({\mathbf{z}})$ with an invertible $\psi$, every task $h:{\mathcal{X}}\to{\mathbb{R}}$ can be equivalently defined via a latent-space function $h^{\prime}:{\mathcal{Z}}\to{\mathbb{R}}$ as $h=h^{\prime}\circ\psi^{-1}$. The space of all possible tasks can thus be represented as
@@
Theorem 4.4 can be viewed as a “no free lunch”-like theorem for representation learning. The original no free lunch theorem (Wolpert, 1996) states that every learner’s performance is equally good when averaged over a uniform distribution on learning problems; here we show that every viable representation is equally complex when averaged over a uniform distribution on tasks. As we will show in Section E.4, the technical intuition of this result is that every representation renders some tasks in the task space “simple” and others “complex”, with the overall task-averaged complexity independent of the particular choice of the representation.
@@
  • A limitation of Theorem 4.8 is that we requires a non-zero probability of explicitly sampling degree-$1$ tasks, in which latent variables are also task outputs. However, we emphasize that this probability is exponentially small as $d$ becomes large, and we conjecture that it can be completely removed in many settings. See Section E.6 for more discussion.
  • Technically, in our proof we introduce a degree analysis of $k$-degree Boolean functions composed with invertible transforms (Lemma E.10), which may be of independent interest in analyzing similar problems.
Connection to the linear representation hypothesis. A number of recent mechanistic interpretation studies show that LLMs often represent abstract, interpretable features as directions in their intermediate representation space (Nanda et al., 2023; Marks & Tegmark, 2023; Gurnee & Tegmark, 2024). Theorem 4.8 could be viewed as a provable, Boolean version of the emergence of such linear representations: permutations and negations are precisely all degree-$1$ Boolean functions, a natural counterpart of degree-$1$ real polynomials (i.e., linear functions) in the real domain. The main significance of this result is that even when proxy tasks involve complex, non-linear functions over true latent variables ${\bm{z}}$, we can still recover ${\bm{z}}$ up to very simple transforms despite the presence of such nonlinearity.
@@ 4.3 Benefits of Learning World Models
Up to now, we have presented sufficient conditions for learning world models. As a complement to these results, this section demonstrates provable benefits of learning world models in the context of an out-of-distribution generalization setting introduced by Abbe et al. (2023).
@@
Let the latent variables during training be uniformly sampled from the Hamming ball $B_{r}\vcentcolon=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$ with $r<d$, and let the latent variables in test be uniformly sampled from ${\mathcal{Z}}$. Let $h:\{\pm 1\}^{m}\to\{\pm 1\}$ be a downstream task such that $h\circ\psi$ is a parity function with $\mathrm{deg}(h)=q>k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose i}\rceil$. Then, if $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and $\mathrm{deg}(h\circ\psi)>k-q+r$, the following holds: (i) the test mean square error (MSE) of any $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$ is larger than $1$; (ii) let $\Phi^{*}$ be a representation that learns the world model up to negations and permutations as in Theorem 4.8 and let $g^{*}$ be a function such that $g^{*}\circ\Phi^{*}\in{\mathcal{H}}(h)$, then the test MSE of $g^{*}\circ\Phi^{*}$ is 00.
@@
As also noted by Abbe et al. (2023), a practical scenario reflected by sampling from $B_{r}$ is length generalization (Anil et al., 2022; Press et al., 2022). Here we show that in this setting, a hierarchical realization with the world model is provably more generalizable than any flat realization, despite that both of them achieve zero i.i.d. test error.
@@ 4.4 Impact of Model Architecture
In the above analysis, we study the role of the low-degree bias with the assumption that the task solutions perfectly adhere to it in the function space. Yet, practical training of neural networks is often more involved than this abstraction. Although neural networks are known as universal function approximators (Hornik et al., 1989; Funahashi, 1989), prior work has shown that models with different architectures may represent the same function differently (Raghu et al., 2021). In particular, embedded nonlinearities such as activation functions can steer how functions are represented by neural networks (Xu et al., 2020; Ziyin et al., 2020).
@@
Notably, for a given input dimension $n$, the Fourier-Walsh transform uses parity functions $\chi_{S}$ as a basis of the $2^{n}$-dimensional vector space ${\mathcal{F}}^{n}=\{f:\{\pm 1\}\to{\mathbb{R}}\}$ (cf. Definition 3.1). To obtain a different basis, we define a basis transform $U$, i.e., an invertible linear transform on ${\mathcal{F}}^{n}$ such that $\{U(\chi_{S})\mid S\subseteq[n]\}$ is a basis of ${\mathcal{F}}^{n}$. The degree of a function $f:\{\pm 1\}^{n}\to{\mathbb{R}}$ under the new basis is then given by
@@ 5 Algorithmic Implications
In this section, we illustrate the algorithmic implications of basis compatibility through two representative tasks: polynomial extrapolation (Xu et al., 2021) and learning physical laws (Kang et al., 2025; Motamed et al., 2025). The first task provides an illustrative example of our results in Section 4.4, while the second is of greater practical interest for learning real-world models and is related to recent efforts in developing video prediction models as world simulators (Brooks et al., 2024). While our experiments are currently limited in scale and mainly serve as proof-of-concept demonstrations, we view extending the ideas discussed in this section to broader contexts as a promising future direction.
@@ 5.1 Polynomial Extrapolation
We begin by a synthetic task in which we train multilayer perceptrons (MLPs) to fit real polynomials $P_{n}(x)=\sum_{i=0}^{n}a_{i}x^{n}$ (see Section F.1 for details). The space of polynomials has a standard basis $\{1,x,x^{2},\ldots\}$, making it a natural counterpart of the Boolean function space with the parity basis. Recovering all basis functions used by $P_{n}$ could generate any polynomial by linearly combining these basis functions, which enables extrapolation.
@@
Motivated by this explanation, here we provide a simple fix: replacing a portion of the ReLU activation functions with functions that are more compatible with the polynomial basis. In practice, we replace half of ReLU functions in each MLP layer by one of the two functions including the identity function $\sigma(x)=x$ and the quadratic function $\sigma(x)=x^{2}$. Representing $f(x)=x^{k}$ would be much easiler using these functions, and we thus expect that a neural network with the low-complexity bias would then use them instead of the remaining ReLU despite the same expressive ability they have, resulting in $\hat{f}\approx f$. As shown in Figure 2(b), this simple method indeed leads to significant improvement in extrapolation. Please see Section G.1 for more results and discussion.
@@
We now show that the approach in Section 5.1 also benefits a sequence prediction task aiming at learning physical laws. Correctly abstracting fundamental physical laws is essential for any model to be a real “world model” for the physical world (Motamed et al., 2025). Notably, a model that “understands” the physical laws is expected to generalize these laws to unseen distributions rather than fit them only in the training domain. Following Kang et al. (2025), we generate two types of object motion sequences reflecting basic physical laws: (i) single-object parabolic motion, and (ii) two-object elastic collision motion (see Section F.2 for more details). We train a transformer (Vaswani et al., 2017) to predict the motion of objects conditional on the first few frames in the motion sequence. The model is then evaluated in an out-of-distribution generalization setup with objects having different initial velocities and sizes. To apply our method, we simply replace every MLP in the transformer with our modified MLP in Section 5.1 (we also replace the remaining ReLUs with GELUs).
As shown in Figure 2(c), our model achieves lower prediction error than transformer in both settings. Note that both models achieve near zero training error. Thus, the fact that our model generalizes better out-of-distribution is not because it fits the data better, but due to it better capturing the underlying laws in object movements. See Section G.2 for the predicted motions from both models.